what is the value of p in 8-3p=13-5 and how did you get that answer?
Found 2 solutions by uma, funmath:
Answer by uma(370)

what is the value of p in 8-3p=13-5 and how did you get that answer?


SOLUTION:
8- 3P = 13 - 5
8 - 3P = 8 ( SUBTRACTING 5 FROM 13)
SUBTRACT 8 FROM BOTH SIDES OF THE EQUATION
==> - 3P = 0
DIVIDE BOTH SIDES OF THE EQUATION BY -3
==> (-3P)/-3 = 0/-3
==> P = 0
VALUE OF P = 0

BYE
PRASEENA.

Answer by funmath(2933)

what is the value of p in 8-3p=13-5 and how did you get that answer?
8-3p=8
-8+8-3p=8-8
-3p=0
-3p/-3=0/-3
p=0
:
Check:
8-3(0)=13-5
8-0=8
8=8 We're right!!!
Happy Calculating!!!!